# Sky-Drive: A Distributed Multi-Agent Simulation Platform for Socially-Aware and Human-AI Collaborative Future Transportation

### v1

@@ Abstract
Recent advances in autonomous system simulation platforms have significantly enhanced the safe and scalable testing of driving policies. However, existing simulators do not yet fully meet the needs of future transportation research—particularly in modeling socially-aware driving agents and enabling effective human-AI collaboration. This paper introduces Sky-Drive, a novel distributed multi-agent simulation platform that addresses these limitations through four key innovations: (a) a distributed architecture for synchronized simulation across multiple terminals; (b) a multi-modal human-in-the-loop framework integrating diverse sensors to collect rich behavioral data; (c) a human-AI collaboration mechanism supporting continuous and adaptive knowledge exchange; and (d) a digital twin (DT) framework for constructing high-fidelity virtual replicas of real-world transportation environments. Sky-Drive supports diverse applications such as autonomous vehicle (AV)–vulnerable road user (VRU) interaction modeling, human-in-the-loop training, socially-aware reinforcement learning, personalized driving policy, and customized scenario generation. Future extensions will incorporate foundation models for context-aware decision support and hardware-in-the-loop (HIL) testing for real-world validation. By bridging scenario generation, data collection, algorithm training, and hardware integration, Sky-Drive has the potential to become a foundational platform for the next generation of socially-aware and human-centered autonomous transportation research. The demo video and code are available at: https://sky-lab-uw.github.io/Sky-Drive-website/.
@@ I Introduction
Autonomous systems and related technologies have made significant strides in recent years, demonstrating increasing maturity in perception, decision-making, and control capabilities [1, 2, 3, 4]. As these technologies continue to advance, future transportation systems are expected to consist of diverse intelligent agents, including autonomous vehicles (AVs), human-driven vehicles (HVs), delivery robots, flying vehicles, and smart infrastructure [5]. In this emerging ecosystem, each agent must not only ensure its own safe and efficient operation, but also align its behavior with human preferences and societal norms through continuous interactions with various road users, such as pedestrians and cyclists. As a result, future research must move beyond the isolated validation of AV performance toward the broader investigation of social awareness and human-AI collaboration in mixed traffic environments.
Validating autonomous driving technologies in real-world settings presents considerable challenges due to safety risks, limited controllability, and the scale of testing required to demonstrate reliability [6, 7, 8]. To mitigate these barriers, the autonomous driving community has developed a variety of simulation platforms, including CARLA [9], AirSim [10], SUMO [11], Vissim [12], Highway-Env [13], MetaDrive [14], SMARTS [15], CarSim [16] and IPG CarMaker [17]. These platforms have significantly accelerated development by providing controlled testing environments. However, they face key limitations in addressing the evolving needs of future transportation research.
First, while existing simulation platforms can emulate multiple agents on a single machine using rule-based or pre-trained learning-based methods [11, 12, 13], they generally do not support real-time participation of human users (e.g., drivers or pedestrians) across multiple terminals. This limitation hinders the collection of authentic human behavior and human-agent interaction data. Such data is particularly valuable for studying rare but safety-critical scenarios—for example, interactions between AVs, HVs, and pedestrians—which pose significant safety and ethical risks when collected in the real world. A distributed simulation platform that enables participants to assume diverse roles across multiple terminals is urgently needed to safely collect such interaction data and to evaluate interaction algorithms in controlled environments.
@@
Figure 1: Overview of Sky-Drive’s key components and functionalities. (a) a distributed multi-agent architecture enabling synchronized simulation across multiple terminals; (b) a multi-modal human-in-the-loop framework capturing comprehensive behavioral data through integrated sensor systems; (c) a digital twin framework that creates high-fidelity virtual replicas of transportation systems through multi-source data integration; (d) a human-AI collaboration mechanism facilitating knowledge exchange between humans and AI systems; (e) the planned integration of foundation models to enhance decision-making, enabling more adaptive and context-aware human-AI collaboration; (f) a hardware-in-the-loop framework, planned for future integration, ensuring that algorithms are evaluated in real-world environments.
Third, although some simulation platforms have integrated reinforcement learning (RL) capabilities to improve autonomous driving policies [14, 15, 28], they remain largely focused on optimizing vehicle-level metrics such as safety, efficiency, and route completion. However, advancing real-world deployment requires moving beyond individual vehicle performance to incorporate social awareness into the decision-making process. Social awareness refers to an autonomous system’s ability to consider the impact of its actions on surrounding road users and the broader traffic environment [29]. This includes promoting traffic flow stability, enhancing the comfort of other participants, and enabling harmonious interactions between AVs and humans in mixed traffic settings. In this context, transportation science offers a valuable foundation. Decades of research have produced validated traffic flow theories and human behavioral models—such as the intelligent driver model (IDM) [30] and the minimizing overall braking induced by lane changes (MOBIL) [31] model—that can inform the design of socially aware autonomous systems.
@@
  • Sky-Drive introduces a distributed multi-agent architecture that enables synchronized simulation across multiple devices through a remote procedure call (RPC) networking model. This design allows independent control of agents on separate terminals while maintaining shared environmental states, better reflecting future mixed traffic.
@@
  • Sky-Drive will integrate foundation models at both the system and agent levels. At the system level, foundation models will provide global observation and feedback to optimize simulation dynamics. At the agent level, they will enhance situational understanding and enable safer, more socially aware, and personalized decision-making.
  • Sky-Drive will incorporate a hardware-in-the-loop (HIL) framework via robot operating system (ROS) integration, enabling direct validation of autonomous driving algorithms on physical vehicles and safe evaluation of algorithms without exposing users to real-world risks.
@@ II-B Human–AI Collaboration Environments
Several simulation platforms have contributed to advancing human-AI collaboration in autonomous driving. For instance, NVIDIA’s DRIVE Sim and Omniverse platform [46] support collaboration by generating physics-based synthetic data for training autonomous systems. However, their approach largely enables one-way knowledge transfer—where simulated data informs AI models—without supporting real-time, bidirectional human-AI interaction. Applied Intuition incorporates human-in-the-loop testing to allow operators to validate autonomous decisions, yet its framework is primarily tailored for offline validation rather than continuous learning [47]. MORAI provides digital twin environments that visualize AI decision-making for human drivers, but its interaction remains limited to basic feedback collection without mechanisms for mutual adaptation or learning [48].
More specialized platforms have made progress toward collaborative learning. MIT’s VISTA enables domain adaptation between real and virtual environments, but focuses mainly on perception rather than interactive decision-making [49]. The GAMMA framework introduces mixed-reality traffic incorporating human behavior, though it lacks explicit mechanisms for integrating human expertise into AI learning [50]. Wayve’s LINGO architecture enhances transparency by providing natural language explanations for AI decisions [51], while SafeMod leverages LLMs for bidirectional planning, mimicking human reasoning in autonomous decision-making [52]. Similarly, SurrealDriver generates realistic driving behaviors that align with human expectations using LLMs [53], and DarwinAI’s GenSynth facilitates collaboration between human designers and AI in developing neural networks for driving tasks [54].
Despite these advances, most platforms still fall short in enabling true human-AI knowledge exchange. They often lack mechanisms for the continuous integration of human feedback, resulting in open-loop rather than closed-loop learning processes. Moreover, few platforms support comprehensive multimodal sensing from humans—such as gaze, voice, physiological signals, and control inputs—which are critical for modeling and understanding driving behaviors. Sky-Drive addresses these limitations through its HAIM and AIHM modules, its multi-modal human-in-the-loop framework, and its closed-loop learning architecture that continuously integrate human expertise into AI development.
@@ III-A Overview
The workflow begins with the DT framework, which creates high-fidelity virtual replicas of transportation systems through multi-source data integration. These virtual environments feed into the distributed multi-agent architecture, enabling synchronized simulations across multiple devices and supporting complex interactions between autonomous agents. The simulation environment created by these two modules serves as the testing ground for the multi-modal human-in-the-loop framework, which captures comprehensive behavioral data from human participants. This data is then processed and utilized by the human-AI collaboration mechanism, facilitating knowledge exchange between humans and autonomous systems. The foundation models integration will enhance system and agent-level capabilities, enabling global observations for performance feedback and aiding individual agents in better understanding human behavior patterns. Finally, the HIL framework connects with the DT framework, enabling real-world algorithm validation while feeding real-world performance data back into the simulation.
@@ III-B1 Scenario Generation & Data Collection
As shown in Fig. 2 (a), this stage employs two complementary approaches to ensure comprehensive scenario coverage: (i) Sky-Drive leverages CARLA and Unreal Engine to generate customizable urban environments with detailed road networks, traffic rules, and environmental conditions, enabling controlled testing of specific driving scenarios. (ii) The DT framework imports real-world data through multi-source integration, including high-precision maps collected by lab-developed AVs, open-source data, and real-world traffic data collection. This data undergoes sophisticated categorization and twinning processes to create digital replicas of physical environments.
@@ III-B3 Hardware Integration & Testing
As illustrated in Fig. 2 (c), this stage bridges simulation and physical deployment through two key components: (i) While the full HIL framework is planned for future development, the current architecture already supports connections to external hardware through standardized ROS interfaces. The lab-developed Ford E-transit electric van serves as the primary testbed, equipped with dashboard monitors, power modules, and a computing rack for algorithm deployment. Testing is primarily conducted at the Madison College Public Safety Training Center in Columbus, WI, and MGA Research Corporation in Burlington, WI. (ii) Sky-Drive also supports testing of vehicle-to-everything (V2X) communication protocols using Cohda Wireless devices, enabling the evaluation of cooperative perception and decision-making across multiple vehicles and infrastructure elements. This testing is crucial for validating the performance of autonomous systems in complex traffic environments.
@@ III-C Case Demonstration
To demonstrate Sky-Drive’s workflow, consider the case of personalized autonomous driving. In this case, Sky-Drive develops an autonomous system that tailors its behavior to the driver’s unique preferences, learning from their driving style and comfort levels.
The workflow begins with the DT framework, which creates high-fidelity virtual replicas of real-world traffic environments using data such as high-precision maps. These environments are then input into the distributed multi-agent architecture, enabling simulations of complex interactions between AVs and humans. During the simulation and training stage, real-time feedback from the driver, such as “It’s too fast” or “The acceleration is too harsh,” is processed by LLMs to infer preferences regarding acceleration and driving style. This feedback is integrated into the human-AI collaboration mechanism, forming a continuous learning loop where the system adapts its driving strategies and provides more personalized guidance. The HIL framework connects the system to physical platforms, validating the personalized driving algorithm in real-world scenarios. This closed-loop workflow enables the development and validation of personalized autonomous systems, from concept testing to real-world deployment, ensuring safety and reliability.
@@ IV Sky-Drive Features
Figure 3: Illustration of Sky-Drive’s distributed multi-agent architecture. Sky-Drive enables synchronized simulation across multiple terminals while maintaining precise real-time interactions between AVs, HVs, and pedestrians through a sophisticated RPC networking model and Socket.IO-based communication platform, supporting comprehensive data collection and real-time analysis of multi-agent behaviors.
@@ IV-A1 System Architecture
At the core of Sky-Drive lies a sophisticated RPC networking model built upon CARLA, using the rpclib library. This extension enhances CARLA’s vehicle control system, enabling crucial improvements for distributed multi-agent simulation. As shown in Fig. 3 (c), Terminal 1 functions as the host (server) that maintains the global simulation environment, while Terminals 2-4 act as clients controlling different agent types. Each terminal independently controls its corresponding agent through various input devices, while ensuring seamless interaction with other agents in the shared environment. The host terminal is responsible for scene customization and map generation, which is then distributed to the client terminals. The scene generation component (Fig. 3 (a)) creates detailed virtual environments with customizable traffic conditions, weather patterns, and road infrastructure, supporting multiple agent types, including AI-controlled AVs, HVs, pedestrians controlled via VR, and rule-based AVs following predefined behaviors (Fig. 1 (a)).
@@ IV-A2 Communication Infrastructure
IV-A3 Real-time Monitoring Platform
A key component of Sky-Drive’s distributed architecture is its real-time monitoring and data management system. Complementing the core networking infrastructure, Sky-Drive has developed a Socket.IO-based communication platform that tracks agent data, including position coordinates, velocity metrics, live video feeds, and sensor readings. As shown in Fig. 3 (b), the platform features a web-based system that provides real-time visualization of agent activities. It streams data to a centralized system where agent interactions are monitored and analyzed in real-time. All simulation data, including agent states, environmental conditions, and interaction events, are logged in a centralized database, enabling comprehensive post-simulation analysis and scenario reproduction.
@@ IV-D Digital Twin Framework
V Sky-Drive Application Case
Figure 4: VR-based experimental setup for studying VRU-AV interactions at unsignalized intersections4 .
V-A VR-based VRU-AV Interaction
Studying interactions between vulnerable road users (VRUs) and AVs is critical for the safe deployment of AV technology in complex urban environments. Although VRU-AV conflicts can lead to serious outcomes, real-world crash data involving these cases remain scarce. More importantly, collecting such data in real traffic is unsafe, difficult to reproduce, and often restricted by ethical constraints. To address this challenge, Sky-Drive provides a VR-enabled platform for investigating VRU-AV interactions in a controlled, immersive, and data-rich environment. Its distributed multi-agent simulation architecture enables synchronized control of multiple agents—across separate terminals and devices—while maintaining real-time coordination. This setup is particularly valuable for modeling high-risk scenarios that are difficult to observe or replicate in the physical world.
As shown in Fig. 4, we conducted a case study focused on right-turn conflicts at unsignalized intersections—a scenario frequently associated with accidents in urban environments. This study leveraged Sky-Drive’s synchronized multi-terminal architecture in a novel experimental setup where human participants experienced the scenario from the pedestrian’s perspective through immersive VR, while researchers controlled an AV making right turns from a separate terminal. During each interaction, Sky-Drive captured multimodal behavioral data from both the AV and the pedestrian. The VR recorded 3D gaze vectors, eye fixations, and reaction times from the pedestrian, while simultaneously logging control signals, deceleration profiles, and trajectory predictions from the AV.
This configuration allows researchers to analyze both the physical outcomes (e.g., successful yielding, near-misses, pedestrian hesitation) and the cognitive-emotional states of the human participant, offering insight into how VRUs perceive and respond to AV behavior.

### v2

# Sky-Drive: A Distributed Multi-Agent Simulation Platform for Human-AI Collaborative and Socially-Aware Future Transportation

@@ Abstract
Recent advances in autonomous system simulation platforms have significantly enhanced the safe and scalable testing of driving policies. However, existing simulators do not yet fully meet the needs of future transportation research—particularly in enabling effective human-AI collaboration and modeling socially-aware driving agents. This paper introduces Sky-Drive, a novel distributed multi-agent simulation platform that addresses these limitations through four key innovations: (a) a distributed architecture for synchronized simulation across multiple terminals; (b) a multi-modal human-in-the-loop framework integrating diverse sensors to collect rich behavioral data; (c) a human-AI collaboration mechanism supporting continuous and adaptive knowledge exchange; and (d) a digital twin framework for constructing high-fidelity virtual replicas of real-world transportation environments. Sky-Drive supports diverse applications such as autonomous vehicle–human road users interaction modeling, human-in-the-loop training, socially-aware reinforcement learning, personalized driving development, and customized scenario generation. Future extensions will incorporate foundation models for context-aware decision support and hardware-in-the-loop testing for real-world validation. By bridging scenario generation, data collection, algorithm training, and hardware integration, Sky-Drive has the potential to become a foundational platform for the next generation of human-centered and socially-aware autonomous transportation systems research. The demo video and code are available at: https://sky-lab-uw.github.io/Sky-Drive-website/.
@@ I Introduction
Autonomous systems and related technologies have made significant strides in recent years, demonstrating increasing maturity in perception, decision-making, and control capabilities [1, 2, 3, 4]. As these technologies continue to advance, future transportation systems are expected to consist of diverse AI-driven intelligent agents, including autonomous vehicles (AVs), human-driven vehicles (HVs), delivery robots, flying vehicles, and smart infrastructure—each capable of perceiving the environment, making decisions, and interacting with others in real time [5]. For these agents to seamlessly integrate into traffic systems governed by humans, they must not only ensure its own safe and efficient operation, but also exhibit acute perceptiveness in recognizing human intentions and preferences (e.g., comfort-oriented speed profiles), while skillfully demonstrating social behaviors comparable to their human counterparts (e.g., yielding behaviors and adherence to implicit right-of-way conventions) [6]. This dual requirement—to understand human expectations and to coordinate with other human road users (HRUs), such as human drivers, pedestrians, and cyclists, to optimize overall system performance—constitutes what we refer to as human-AI collaboration and social awareness, respectively. Therefore, future research should transcend the isolated validation of AV safety performance toward a more comprehensive investigation of human-AI collaboration and social awareness in mixed traffic environments, where multiple types of HRUs with varying levels of autonomy interact continuously.
Validating autonomous driving technologies in real-world settings presents considerable challenges due to safety risks, limited controllability, and the scale of testing required to demonstrate reliability [7, 8, 9]. To mitigate these barriers, the autonomous driving community has developed a variety of simulation platforms, including CARLA [10], AirSim [11], SUMO [12], Vissim [13], Highway-Env [14], MetaDrive [15], SMARTS [16], CarSim [17] and IPG CarMaker [18]. These platforms have significantly accelerated development by providing controlled testing environments. However, they face key limitations in addressing the evolving needs of future transportation research. First, while existing simulation platforms can emulate multiple agents on a single machine using rule-based or pre-trained learning-based methods [12, 13, 14], they generally do not support real-time participation of HRUs across multiple terminals. This limitation hinders the collection of authentic human behavior and HRUs’ interaction data. Such data is particularly valuable for studying rare but safety-critical scenarios—for example, interactions between AVs, HVs, and pedestrians—which pose significant safety and ethical risks when collected in the real world. A distributed simulation platform that enables participants to assume diverse roles across multiple terminals is urgently needed to safely collect such interaction data and to evaluate interaction algorithms in controlled environments.
@@
Figure 1: Overview of Sky-Drive’s key components and functionalities. (a) a distributed multi-agent architecture enabling synchronized simulation across multiple terminals; (b) a multi-modal human-in-the-loop framework capturing comprehensive behavioral data through integrated sensor systems; (c) a digital twin framework that creates high-fidelity virtual replicas of transportation systems through multi-source data integration; (d) a human-AI collaboration mechanism facilitating knowledge exchange between humans and AI systems; (e) the planned integration of foundation models to enhance decision-making, enabling more adaptive and context-aware human-AI collaboration; (f) a hardware-in-the-loop framework, planned for future integration, enabling remote control and data collection and ensuring that algorithms are evaluated in real-world environments.
Third, although some simulation platforms have integrated reinforcement learning (RL) capabilities to improve autonomous driving policies [15, 16, 29], they remain largely focused on optimizing vehicle-level metrics such as safety, efficiency, and route completion. However, advancing real-world deployment requires moving beyond individual vehicle performance to incorporate social awareness into the decision-making process. Social awareness refers to an autonomous system’s ability to coordinate with other traffic participants to enhance overall system performance [30]. This includes promoting traffic flow stability, mitigating congestion through cooperative maneuvers, enhancing the collective comfort of all road users, and enabling harmonious interactions between AVs and humans in mixed traffic settings. In this context, transportation science offers a valuable foundation. Decades of research have produced validated traffic flow theories and human behavioral models—such as the intelligent driver model (IDM) [31] and the minimizing overall braking induced by lane changes (MOBIL) [32] model—that can inform the design of socially aware autonomous systems capable of optimizing not just individual performance but the efficiency and safety of the entire traffic ecosystem.
@@
  • Sky-Drive introduces a distributed multi-agent architecture that enables synchronized simulation across multiple devices through remote procedure call (RPC). This design allows independent control of agents on separate terminals while maintaining shared environmental states, better reflecting future mixed traffic.
@@
  • Sky-Drive will integrate foundation models, i.e., LLM and VLM, at both the system and agent levels. At the system level, foundation models will provide global observation and feedback to optimize simulation dynamics. At the agent level, they will enhance situational understanding and enable safer, more socially aware, and personalized decision-making.
  • Sky-Drive will incorporate a hardware-in-the-loop (HIL) framework via robot operating system (ROS), enabling direct validation of autonomous driving algorithms on physical vehicles and safe evaluation of algorithms without exposing users to real-world risks.
@@ II-B Human–AI Collaboration Environments
Several simulation platforms have contributed to advancing human-AI collaboration in autonomous driving. For instance, NVIDIA’s DRIVE Sim and Omniverse platform [47] support collaboration by generating physics-based synthetic data for training autonomous systems. However, their approach largely enables one-way knowledge transfer—where simulated data informs AI models—without supporting real-time, bidirectional human-AI interaction. Applied Intuition [48] incorporates human-in-the-loop testing to allow operators to validate autonomous decisions, yet its framework is primarily tailored for offline validation rather than continuous learning. MORAI [49] provides digital twin environments that visualize AI decision-making for human drivers, but its interaction remains limited to basic feedback collection without mechanisms for mutual adaptation or learning.
More specialized platforms have made progress toward collaborative learning. MIT’s VISTA [50] enables domain adaptation between real and virtual environments, but focuses mainly on perception rather than interactive decision-making. The GAMMA framework [51] introduces mixed-reality traffic incorporating human behavior, though it lacks explicit mechanisms for integrating human experience into AI learning. Wayve’s LINGO architecture [52] enhances transparency by providing natural language explanations for AI decisions, while SafeMod [53] leverages LLMs for bidirectional planning, mimicking human reasoning in autonomous decision-making. Similarly, SurrealDriver [54] generates realistic driving behaviors that align with human expectations using LLMs, and DarwinAI’s GenSynth [55] facilitates collaboration between human designers and AI in developing neural networks for driving tasks.
Despite these advances, most platforms still fall short in enabling true human-AI knowledge exchange. They often lack mechanisms for the continuous integration of human feedback, resulting in open-loop rather than closed-loop learning processes. Moreover, few platforms support comprehensive multimodal sensing from humans—such as gaze, voice, physiological signals, and control inputs—which are critical for modeling and understanding driving behaviors. Sky-Drive addresses these limitations through its HAIM and AIHM modules, its multi-modal human-in-the-loop framework, and its closed-loop learning architecture that continuously integrates human experience into AI development.
@@ III-A Overview
As shown in Fig. 1, the workflow begins with ❶ the DT framework, which creates high-fidelity virtual replicas of transportation systems through multi-source data integration. These virtual environments feed into ❷ the distributed multi-agent architecture, enabling synchronized simulations across multiple devices and supporting complex interactions between autonomous agents. Together, the DT framework and the multi-agent architecture form the simulation environment that serves as the testing ground for ❸ the multi-modal human-in-the-loop framework, which captures comprehensive behavioral data from human participants. ❹ The human-AI collaboration mechanism then processes and utilizes this data, facilitating knowledge exchange between humans and autonomous systems. ❺ The foundation models integration will enhance system and agent-level capabilities, enabling observations for performance feedback and aiding individual agents in better understanding human behavior patterns. Finally, ❻ the HIL framework connects with the DT framework, enabling real-world algorithm validation while feeding real-world performance data back into the simulation.
@@ III-B1 Scenario Generation & Data Collection
As shown in Fig. 2 (a), this stage employs two complementary approaches to ensure comprehensive scenario coverage: (i) Sky-Drive leverages CARLA and Unreal Engine to generate customizable urban environments with detailed road networks, traffic rules, and environmental conditions, enabling controlled testing of specific driving scenarios. (ii) The DT framework imports real-world data through multi-source integration, including high-definition maps collected by lab-developed AVs, open-source data, and real-world traffic data collection. This data undergoes categorization and twinning processes to create digital replicas of physical environments.
@@ III-B3 Hardware Integration & Testing
As illustrated in Fig. 2 (c), this stage bridges simulation and physical deployment through two key components: (i) While the full HIL framework is planned for future development, the current architecture already supports connections to external hardware through standardized ROS. The lab-developed Ford E-transit electric van serves as the primary testbed, equipped with dashboard monitors, power modules, drive-by-wire systems, and a computing rack for algorithm deployment. Testing is primarily conducted at the Madison College Public Safety Training Center in Columbus, WI, and MGA Research Corporation in Burlington, WI. (ii) Sky-Drive also integrates a 5G mmWave mobile hotspot to support low-latency teleoperation, enabling remote human operators to monitor and, when necessary, take control of physical vehicles in real time.
@@ III-C Case Demonstration
To demonstrate Sky-Drive’s workflow, we consider the case of personalized autonomous driving. In this case, Sky-Drive develops an autonomous system that tailors its behavior to the driver’s unique preferences, learning from their driving styles and comfort levels.
The workflow begins with the DT framework, which creates high-fidelity virtual replicas of real-world traffic environments using data such as high-definition maps. These environments are then input into the distributed multi-agent architecture, enabling simulations of complex interactions between AVs and humans. During the simulation and training stage, real-time feedback from the driver, such as “It’s too fast” or “The acceleration is too harsh,” is processed by LLMs to infer preferences regarding acceleration and driving style. This feedback is integrated into the human-AI collaboration mechanism, forming a continuous learning loop where the system adapts its driving strategies and provides more personalized guidance. The HIL framework connects the system to physical platforms, validating the personalized driving algorithm in real-world scenarios. This closed-loop workflow enables the development and validation of personalized autonomous systems, from concept testing to real-world deployment, ensuring safety and reliability.
@@ IV Sky-Drive Features
Figure 3: Illustration of Sky-Drive’s distributed multi-agent architecture. Sky-Drive enables synchronized simulation across multiple terminals while maintaining precise real-time interactions between AVs, HVs, and pedestrians through RPC and Socket.IO-based communication platform, supporting comprehensive data collection and real-time analysis of multi-agent behaviors.
@@ IV-A1 System Architecture
At the core of Sky-Drive lies RPC built upon CARLA, using the rpclib library. This extension enhances CARLA’s vehicle control system, enabling crucial improvements for distributed multi-agent simulation. As shown in Fig. 3 (c), Terminal 1 functions as the host (server) that maintains the global simulation environment, while Terminals 2-4 act as clients controlling different agent types. Each terminal independently controls its corresponding agent through various input devices, while ensuring seamless interaction with other agents in the shared environment. The host terminal is responsible for scene customization and map generation, which is then distributed to the client terminals. The scene generation component (Fig. 3 (a)) creates detailed virtual environments with customizable traffic conditions, weather patterns, and road infrastructure, supporting multiple agent types, including AI-controlled AVs, HVs, pedestrians controlled via VR, and rule-based AVs following predefined behaviors (Fig. 1 (a)).
@@ IV-A2 Communication Infrastructure
IV-A3 Real-time Monitoring Webstie
A key component of Sky-Drive’s distributed architecture is its real-time monitoring and data management website. Complementing the core communication infrastructure, Sky-Drive has developed a Socket.IO-based communication platform that tracks agent data, including position coordinates, velocity metrics, live video feeds, and sensor readings. As shown in Fig. 3 (b), the platform features a web-based system that provides real-time visualization of agent activities. It streams data to a centralized system where agent interactions are monitored and analyzed in real-time. All simulation data, including agent states, environmental conditions, and interaction events, are logged in a centralized database, enabling comprehensive post-simulation analysis and scenario reproduction.
@@ IV-D Digital Twin Framework
V Sky-Drive Use Case
Figure 4: VR-based experimental setup for studying AV-HRU interactions at unsignalized intersections.
V-A VR-based AV-HRU Interaction
Studying interactions between AVs and HRUs is critical for the safe deployment of AV technology in complex urban environments. Although there are reported incidents of AV-HRU conflicts, real-world crash data involving such cases remains scarce. More importantly, collecting such data in real traffic is unsafe, difficult to reproduce, and often restricted by ethical constraints. To address this challenge, Sky-Drive provides a VR-enabled platform for investigating AV-HRU interactions in a controlled, immersive, and data-rich environment. It’s a distributed multi-agent simulation architecture that enables synchronized control of multiple agents—across separate terminals and devices—while maintaining real-time coordination. This setup is particularly valuable for modeling high-risk scenarios that are difficult to observe or replicate in the physical world.
As shown in Fig. 4, we conducted a case study focused on right-turn conflicts at unsignalized intersections—a scenario frequently associated with accidents in urban environments. This study leveraged Sky-Drive’s synchronized multi-terminal architecture in a novel experimental setup where human participants experienced the scenario from the pedestrian’s perspective through immersive VR, while researchers monitored the decision-making of an AV from a separate terminal and occasionally intervened when necessary. During each interaction, Sky-Drive captured multimodal behavioral data from both the AV and the pedestrian. The VR recorded 3D gaze vectors, eye fixations, and reaction times from the pedestrian, while simultaneously logging control signals, deceleration profiles, and trajectory predictions from the AV.
This configuration allows researchers to analyze both the physical outcomes (e.g., successful yielding, near-misses, pedestrian hesitation) and the cognitive-emotional states of the human participant, offering insight into how HRUs perceive and respond to AV behavior.
TABLE II: The performance of PPO, HACO, and HAIM-DRL methods.
Method Test Safety Violation $\downarrow$ Test Return $\uparrow$ Test Disturbance Rate $\downarrow$ Test Success Rate $\uparrow$ Train Samples $\downarrow$
PPO 80.84 1591.00 - 0.35 500,000
HACO 12.14 1578.43 0.0137 0.35 8,000
HAIM-DRL (ours) 11.25 1590.85 0.0121 0.38 8,000
TABLE III: Performance comparison with baselines during testing. Mean and standard deviation over 3 seeds.
VLM-SR 0.53 $\pm$ 0.27 0.02 $\pm$ 0.00 47.9 $\pm$ 9.2 0.18 $\pm$ 0.25 0.0 $\pm$ 0.0
LORD 0.17 $\pm$ 0.08 0.02 $\pm$ 0.02 45.1 $\pm$ 57.1 0.02 $\pm$ 0.02 0.0 $\pm$ 0.0